# MoFO: Momentum-Filtered Optimizer for Mitigating Forgetting in LLM Fine-Tuning

### v3

@@ 3.2 Initial Analysis on Forgetting Mitigation
Suppose the parameter space is $\mathbb{R}^{d}$, and the updating ratio is $\alpha\%=1/d$, i.e., only one coordinate is updated in each iteration. We assume the pre-training loss is $\mathcal{L}_{\rm pretrain}(\theta)=\frac{1}{2}\|\theta\|_{2}^{2}$ and that the model has been trained to the global minimum $\theta_{\rm pretrain}=(0,0,\dots,0)$ during the pre-training phase. The fine-tuning loss is given by $\mathcal{L}(\theta)=\prod_{i=1}^{d}(a_{i}\theta_{i}-b_{i})^{2}$, where $a_{i},b_{i}>0$ for any $1\leq i\leq d$. In this example, the set of global minima of $\mathcal{L}(\theta)$ is a union of hyperplanes:
$S=\bigcup_{i=1}^{d}S_{i},\quad\text{where }S_{i}:=\{\theta\in\mathbb{R}^{d}:\theta_{i}=b_{i}/a_{i}\}.$
@@
We note that the loss landscape of neural networks are highly non-convex (Liu et al., 2022). Here, we also adopt a non-convex fine-tuning loss $\mathcal{L}(\theta)$. Further, we note that the set of global-minima $S$ consists of infinitely many minima spread across multiple hyperplanes, aligning with the observation on the degenerate structure of minima in neural networks (Lin et al., 2024b).
@@
Figure 3: The fine-tuning loss landscape and the training paths of different optimization methods. A logarithmic scale is applied to the loss values for better visualization. MoFO converges to a minimum closest to the pre-trained model.
@@
In contrast, MoFO is influenced by only one attractor ($\theta_{1}=1$), leading it to follow the green line and converge to $(1,0)$, which achieves a lower pre-training loss of 0.5333By symmetry, MoFO may also be influenced by another attractor $\theta_{2}=1$ and converges vertically to $(0,1)$.. We hypothesize that, for full-parameter fine-tuning (Adam), interference among multiple attractors drives convergence toward a “balanced” solution, whereas MoFO mitigates such interference by selectively updating only parameters with large momentum magnitudes. Such an updating rule helps MoFO move toward a single attractor, thereby converging closer and forgetting less. In addition to Adam and MoFO, we provide convergence paths of some other baseline methods in Appendix C.2.
@@ 4.1 Experimental Settings
We verify the effectiveness of MoFO on instruction fine-tuning and continual fine-tuning. We use Llama-2-7B (Touvron et al., 2023), Gemma-2B-IT (Team et al., 2024), and TinyLlama-1.1B (Zhang et al., 2024b) as our base models. The instruction fine-tuning datasets cover question-answer pairs from different domains like mathematical reasoning and medical knowledge. Specifically, the datasets include: MetaMathQA (Yu et al., 2024a) and PMC-LLaMA-Instructions (Wu et al., 2024). We randomly sample 39.5K and 51K instances from these datasets, respectively, for training the LLMs. Additionally, We investigate the performance of MoFO in the continual fine-tuning scenario by implementing our approach on the TRACE benchmark dataset (Wang et al., 2023b).
@@
For more descriptions and implementation details of these metrics and datasets, see Appendix D.
@@ 4.2 Instruction Fine-Tuning
In this section, we investigate the effectiveness of the MoFO algorithm in both preserving general capabilities and learning fine-tuning tasks. The implementation details are provided in Appendix D. The specific hyperparameter settings in each experiment are provided in Appendix D.3.
@@
We fine-tune the Llama-2-7B model on the MetaMathQA dataset using $L_{1}$ and $L_{2}$ regularization, as well as LoRA, and compare their performance with MoFO. We present the results in Figure LABEL:pareto and plot Pareto optimal fronts [^5] Since it is impractical to exhaust all hyperparameter configurations in real experiments, we present linear interpolation approximations of the Pareto fronts in Figure LABEL:pareto. for these methods. Details of the hyperparameter configurations for this experiment are provided in Appendix D.3. These results show the effectiveness of the MoFO algorithm in both optimization and mitigating forgetting.
The result reveals that MoFO consistently achieves a better Pareto front in comparison to baseline methods. When compared to regularization methods and LoRA, MoFO exhibits less forgetting and can even maintain general capabilities with comparable GSM8K accuracies. Additionally, MoFO outperforms regularization methods in math tasks when the magnitudes of forgetting are similar. We also note that $L_{1}$ and $L_{2}$ regularization (Panigrahi et al., 2023; Li et al., 2018) require storing the pre-training weights throughout the entire fine-tuning process for regularization computation, which incurs additional memory overhead.
Additional experimental results of instruction fine-tuning using other LLMs, datasets, and baselines can be found in Appendix F.
@@
Figure LABEL:distance_chek shows that models fine-tuned with MoFO are closer to the pre-trained model compared to other baseline methods.
@@ 4.3 Continual Fine-Tuning
In this section, we explore the performance of our proposed MoFO in continual fine-tuning on the TRACE benchmark (Wang et al., 2023b). We sequentially train TinyLlama-1.1B on the TRACE dataset, which includes the eight tasks from different domains. The implementation details are provided in Appendix D.
@@ 4.4 Impact of Update Strategy in MoFO
In addition to MoFO, we consider three other BCD methods, random BCD, gradient-filtered BCD, and MV-filtered BCD. Random BCD updates a random subset of parameters at each iteration. Gradient-filtered BCD replaces MoFO’s filter $\texttt{FLT}_{\alpha}(m_{t})$ with $\texttt{FLT}_{\alpha}(g_{t})$, while MV-filtered BCD uses $\texttt{FLT}_{\alpha}(m_{t}/\sqrt{v_{t}})$.
@@
In terms of GSM8K performance, our proposed MoFO method significantly surpasses Random BCD, Gradient-filtered BCD, and MV-filtered BCD, indicating that updating parameters with the largest momentum leads to strong optimization power. More insights towards this result are provided in Appendix LABEL:subapp:insight-update-strategy.
@@ 4.5 Furthur Analysis
Guidelines for setting $\alpha\%$. Experiments show that setting the updating fraction $\alpha\%=15\%$ works the best for most of our experiments; and $5\%-15\%$ all work quite well. We provide a more detailed guideline for determining $\alpha\%$ in Appendix E. The guideline involves randomly sampling a small proxy subset and performing a grid search over possible $\alpha\%$ values.
@@
Efficiency Analysis. We provide an efficiency analysis on MoFO in Appendix LABEL:subapp:efficiency_mofo. The results show that MoFO requires only around 4%–5% additional training time compared with Default FT throughout the entire training process.
@@ 5 Conclusion and Limitations
Impact Statements
We design new algorithms to mitigate the forgetting issues for LLMs. Our work can help the community better train large AI model training. However, it would be a potential threat if the AI models are used for illegal usage.
Appendix A Related Works
A.1 Forgetting in Continual Learning
Catastrophic forgetting, a significant issue where models forget previously learned information upon learning new data, has received considerable attention in machine learning (McCloskey & Cohen, 1989; Goodfellow et al., 2013; Kemker et al., 2018; Ramasesh et al., 2021; Liu et al., 2024). Traditional continual learning primarily focuses on addressing catastrophic forgetting in sequential-task learning scenarios. In addition to investigating the forgetting of pre-training knowledge during fine-tuning, Section 4.3 conducts experimental studies on catastrophic forgetting in sequential-task fine-tuning processes, which aligns more closely with conventional continual learning paradigms.
Regularization-based methods. These methods introduce constraints to the training process to preserve past knowledge, such as adding regularization to the loss functions (Kirkpatrick et al., 2017; Aljundi et al., 2018; Zenke et al., 2017; Li et al., 2018; Ritter et al., 2018; Kumar et al., 2023) or the embedding/output changes (Li & Hoiem, 2017; Rannen et al., 2017; Buzzega et al., 2020; Huang et al., 2021; Cha et al., 2020). Some regularization-based approaches still rely on partial information from previous models (Kirkpatrick et al., 2017). In contrast, MoFO does not require past information and does not alter the original loss function, making it inherently orthogonal to regularization-based methods.
Optimization-based methods. These methods focus on modifying the training algorithm to mitigate forgetting. In traditional continual learning, optimization-based methods commonly include, but are not limited to, gradient projection techniques (Wang et al., 2023a; Lopez-Paz & Ranzato, 2017), meta-learning approaches (Beaulieu et al., 2020; Javed & White, 2019), and strategies leveraging the structure of the loss landscapes (Mirzadeh et al., 2020a, b). When it comes to forgetting-mitigation in LLM training, recent studies have explored optimization strategies that update only a subset of parameters at each iteration. For instance, Hui et al. (2024) randomly freezes half of the model’s parameter modules and updatets the rest at each iteration. Ke et al. (2023b, a) introduce a soft-masking mechanism that selects parameters for update based on their importance values. Further, Zhang et al. (2024a) combines selective module updating with soft-masking. MoFO, which also falls into this category, updates parameters with largest momentum magnitudes at each iteration. Compared to these works, our study provides a theoretical convergence guarantee of our proposed method, thereby establishing its effectiveness in LLM fine-tuning.
Model merging methods. These methods balance learning new knowledge and retaining old knowledge by merging the new and past models. One line of research focuses on model averaging, which interpolates between the weights of different LLMs (Wortsman et al., 2022a, b; Eeckt et al., 2022; Yadav et al., 2024; Lin et al., 2023, 2024a). Another line of research relies on the observation that task-specific knowledge largely resides in a subspace of the weight space (Ilharco et al., 2023; Panigrahi et al., 2023; Gueta et al., 2023; Zhu et al., 2024), and leverage task vectors or task localization to preserve pre-training knowledge in the fine-tuned models (Panigrahi et al., 2023; Yadav et al., 2024; Yu et al., 2024b).
Architecture-based methods. These methods modify the model’s architecture in training. LoRA (Hu et al., 2022), as the most popular parameter-efficient fine-tuning (PEFT) method, freezes the pre-training weights and introduces low-rank trainable matrices. Variants of LoRA are applied in continual learning for LLMs (Ren et al., 2024; Wang et al., 2023a). However, LoRA is observed to forget less but also learn less than default fine-tuning (Biderman et al., 2024). Other approaches adaptively expand model capacity or isolate partial weights to mitigate interference between new and old tasks (Wang et al., 2023a; Razdaibiedina et al., 2023). In contrast, MoFO updates a subset of parameters at each iteration, but does not alter the total trainable parameters.
A.2 Block Coordinate Descent
Block Coordinate Descent (BCD) involves iteratively optimizing over a block of coordinates while holding the others constant. The foundational work of (Tseng, 2001) provides a comprehensive analysis of the convergence properties of BCD under certain conditions. Subsequent research has explored various BCD variants (Hong et al., 2017), including random BCD (Nesterov, 2012; Richtárik & Takáč, 2014; Lu & Xiao, 2015), cyclic BCD (Sun & Hong, 2015), and greedy BCD (Nutini et al., 2015). Among these, the greedy variant, also known as Gauss-Southwell BCD method, has drawn attention due to its ability to prioritize coordinates that yield the most substantial improvement in each iteration, thereby potentially accelerating convergence.
In the realm of machine learning, BCD has also found applications (Nutini et al., 2022). For example, Luo et al. (2024) leverages BCD to perform memory-efficient fine-tuning of LLM and (Xu & Zhang, 2024) uses random masking to perform this. In federated learning, (Rothchild et al., 2020) adopts top-$k$ momentum value unsketch rather than our top-$k$ momentum filtering to tackle communication bottleneck and convergence issues. In LLMs, some concurrent works propose BCD-based algorithms leveraging task vectors to enhance fine-tuning performance (Li et al., 2024) and mitigate catastrophic forgetting in multi-task learning (Panda et al., 2024). Our approach can be regarded as a type of greedy BCD adapted to Adam, achieving good performance in fine-tuning tasks and alleviating forgetting.
Appendix B Theoretical Analysis
B.1 Supplementary Analysis on the Top-$\alpha\%$ Filter

### v4

@@ 3.2 Initial Analysis on Forgetting Mitigation
Suppose the parameter space is $\mathbb{R}^{d}$, and the updating ratio is $\alpha=1/d$, i.e., only one coordinate is updated in each iteration. We assume the pre-training loss is $\mathcal{L}_{\rm pretrain}(\theta)=\frac{1}{2}\|\theta\|_{2}^{2}$ and that the model has been trained to the global minimum $\theta_{\rm pretrain}=(0,0,\dots,0)$ during the pre-training phase. The fine-tuning loss is given by $\mathcal{L}(\theta)=\prod_{i=1}^{d}(a_{i}\theta_{i}-b_{i})^{2}$, where $a_{i},b_{i}>0$ for any $1\leq i\leq d$. In this example, the set of global minima of $\mathcal{L}(\theta)$ is a union of hyperplanes:
$S=\bigcup_{i=1}^{d}S_{i},\quad\text{where }S_{i}=\{\theta\in\mathbb{R}^{d}:\theta_{i}=b_{i}/a_{i}\}.$
@@
We note that the loss landscapes of neural networks are generally non-convex (Liu et al., 2022). Here, we also adopt a non-convex fine-tuning loss $\mathcal{L}(\theta)$. Further, we note that the set of global-minima $S$ consists of infinitely many minima spread across multiple hyperplanes, aligning with the observation on the degenerate structure of minima in neural networks (Lin et al., 2024b).
@@
Figure 3: The fine-tuning loss landscape and the training paths of different optimization methods. The color bar indicates the fine-tuning loss value.
A logarithmic scale is applied to the loss values for better visualization. MoFO converges to a minimum closest to the pre-trained model.
@@
In contrast, MoFO is influenced by only one attractor ($\theta_{1}=1$), leading it to follow the green line and converge to $(1,0)$, which achieves a lower pre-training loss of 0.5444By symmetry, MoFO may also be influenced by another attractor $\theta_{2}=1$ and converges vertically to $(0,1)$.. We hypothesize that, for full-parameter fine-tuning (Adam), interference among multiple attractors drives convergence toward a “balanced” solution, whereas MoFO mitigates such interference by selectively updating only parameters with large momentum magnitudes. Such an updating rule helps MoFO move toward a single attractor, thereby converging closer and forgetting less. In addition to Adam and MoFO, we provide convergence paths of some other baseline methods in Appendix B.4.
@@ 4.1 Experimental Settings
We verify the effectiveness of MoFO on instruction fine-tuning and continual fine-tuning. We use Llama-2-7B (Touvron et al., 2023), Gemma-2B-IT (Team et al., 2024), and TinyLlama-1.1B (Zhang et al., 2024b) as our base models. The instruction fine-tuning datasets cover question-answer pairs from different domains like mathematical reasoning and medical knowledge. Specifically, the datasets include: MetaMathQA (Yu et al., 2024b), PMC-LLaMA-Instructions (Wu et al., 2024), Magicoder-Evol-Instruct (Wei et al., 2023). We randomly sample 39.5K and 51K instances from these datasets, respectively, for training the LLMs. Additionally, We investigate the performance of MoFO in the continual fine-tuning scenario by implementing our approach on the TRACE benchmark dataset (Wang et al., 2023b).
@@
For more descriptions and implementation details of these metrics and datasets, see Appendix C.
@@ 4.2 Instruction Fine-Tuning
In this section, we investigate the effectiveness of the MoFO algorithm in both preserving general capabilities and learning fine-tuning tasks. The implementation details are provided in Appendix C. The specific hyperparameter settings in each experiment are provided in Appendix C.3.
@@
We fine-tune the Llama-2-7B model on the MetaMathQA dataset using $L_{1}$ and $L_{2}$ regularization, as well as LoRA, and compare their performance with MoFO. We present the results in Figure 5(a) and plot Pareto optimal fronts [^6] Since it is impractical to exhaust all hyperparameter configurations in real experiments, we present linear interpolation approximations of the Pareto fronts in Figure 5(a). for these methods. Details of the hyperparameter configurations for this experiment are provided in Appendix C.3. These results show the effectiveness of the MoFO algorithm in both optimization and mitigating forgetting.
(a) The performance on the math task (GSM8K) and the scores in general capabilities of Llama-2-7B after fine-tuning on the MetaMathQA dataset. Only points on the Pareto front are shown as solid points, while the remaining points are presented as semi-transparent. The results show that compared with $L_{1}$, $L_{2}$ regularization, and LoRA across various hyperparameter configurations, the MoFO algorithm achieves a better Pareto front.
The result reveals that MoFO consistently achieves a better Pareto front in comparison to baseline methods. When compared to regularization methods and LoRA, MoFO exhibits less forgetting and can even maintain general capabilities with comparable GSM8K accuracies. Additionally, MoFO outperforms regularization methods in math tasks when the magnitudes of forgetting are similar. We also note that $L_{1}$ and $L_{2}$ regularization (Panigrahi et al., 2023; Li et al., 2018) require storing the pre-training weights throughout the entire fine-tuning process for regularization computation, which incurs additional memory overhead. For preliminary analysis on why MoFO might compare favorably to $L_{1}/L_{2}$ regularization, see Appendix F.5.
Appendix E reports additional instruction fine-tuning results. Specifically, we evaluate
  • LLM variants: Gemma-2B-IT and Llama-2-7B-Chat;
  • Domain-specific datasets: medical dataset (PMC-LLaMA-Instruct (Wu et al., 2024)), coding dataset (Magicoder-Evol-Instruct (Wei et al., 2023));
  • Different baselines: HMA (Lin et al., 2024a), CoFiTune (Zhang et al., 2024a), Soft-masking (Ke et al., 2023a; b).
@@
Figure 5(b) shows that models fine-tuned with MoFO are closer to the pre-trained model compared to other baseline methods.
(b) The distances for the fine-tuned Llama2-7B on MetaMathQA. The results show MoFO achieves minima closer to the pre-trained model.
@@ 4.3 Continual Fine-Tuning
In this section, we explore the performance of our proposed MoFO in continual fine-tuning on the TRACE benchmark (Wang et al., 2023b). We sequentially train TinyLlama-1.1B on the TRACE dataset, which includes the eight tasks from different domains. The implementation details are provided in Appendix C.
@@ 4.4 Impact of Update Strategy in MoFO
In addition to MoFO, we consider three other BCD methods with different filtering strategies: random BCD, gradient-filtered BCD, and MV-filtered BCD. Random BCD updates a random subset of parameters at each iteration. Gradient-filtered BCD replaces MoFO’s filter $\texttt{FLT}_{\alpha}(m_{t})$ with $\texttt{FLT}_{\alpha}(g_{t})$, while MV-filtered BCD uses $\texttt{FLT}_{\alpha}(m_{t}/\sqrt{v_{t}})$.
@@
In terms of GSM8K performance, our proposed MoFO method significantly surpasses Random BCD, Gradient-filtered BCD, and MV-filtered BCD, indicating that updating parameters with the largest momentum leads to strong optimization power. Additional comparative experiments on BCD filtering strategies are presented in Appendix F.3. More insights towards this result are provided in Appendix F.4.
@@ 4.5 Furthur Analysis
Guidelines for setting $\alpha$. Experiments show that setting the updating fraction $\alpha=15\%$ works the best for most of our experiments; and $5\%-15\%$ all work quite well. We provide a more detailed guideline for determining $\alpha$ in Appendix D. The guideline involves randomly sampling a small proxy subset and performing a grid search over possible $\alpha$ values.
@@
Efficiency Analysis. We provide an efficiency analysis on MoFO in Appendix F.2. The results show that MoFO requires only around $4\%-5\%$ additional training time compared with Default FT throughout the entire training process.
5 Related Works
5.1 Forgetting in Continual Learning
Catastrophic forgetting, a significant issue where models forget previously learned information upon learning new data, has received considerable attention in machine learning (McCloskey & Cohen, 1989; Goodfellow et al., 2013; Kemker et al., 2018; Ramasesh et al., 2021; Verwimp et al., 2023; Liu et al., 2024). Traditional continual learning primarily focuses on addressing catastrophic forgetting in sequential-task learning scenarios. In addition to investigating the forgetting of pre-training knowledge during fine-tuning, Section 4.3 conducts experimental studies on catastrophic forgetting in sequential-task fine-tuning processes, which aligns more closely with conventional continual learning paradigms.
Regularization-based methods. These methods introduce constraints to the training process to preserve past knowledge, such as adding regularization to the loss functions (Kirkpatrick et al., 2017; Aljundi et al., 2018; Zenke et al., 2017; Li et al., 2018; Ritter et al., 2018; Kumar et al., 2023) or the embedding/output changes (Li & Hoiem, 2017; Rannen et al., 2017; Buzzega et al., 2020; Huang et al., 2021; Cha et al., 2020). Some regularization-based approaches still rely on partial information from previous models (Kirkpatrick et al., 2017). In contrast, MoFO does not require past information and does not alter the original loss function, making it inherently orthogonal to regularization-based methods. To improve generalization and robustness to noise after instruction-tuning, several studies introduce explicit regularization (Li & Zhang, 2021; Zhang et al., 2023). In particular, Zhang et al. (2023) proposes a Hessian-based penalty that encourages convergence to flatter minima. It is an interesting direction for future research to evaluate whether MoFO’s momentum filtering mechanism implicitly favors more stable minima, thereby further enhancing generalization and noise robustness.
Optimization-based methods. These methods focus on modifying the training algorithm to mitigate forgetting. In traditional continual learning, optimization-based methods commonly include, but are not limited to, gradient projection techniques (Wang et al., 2023a; Lopez-Paz & Ranzato, 2017), meta-learning approaches (Beaulieu et al., 2020; Javed & White, 2019), and strategies leveraging the structure of the loss landscapes (Mirzadeh et al., 2020a; b). When it comes to forgetting-mitigation in LLM training, recent studies have explored optimization strategies that update only a subset of parameters at each iteration. For instance, Hui et al. (2024) randomly freezes half of the model’s parameter modules and updatets the rest at each iteration. Ke et al. (2023b; a) introduce a soft-masking mechanism that selects parameters for update based on their importance values. Further, Zhang et al. (2024a) combines selective module updating with soft-masking. MoFO, which also falls into this category, updates parameters with largest momentum magnitudes at each iteration. Compared to these works, our study provides a theoretical convergence guarantee of our proposed method, thereby establishing its effectiveness in LLM fine-tuning.
Model merging methods. These methods balance learning new knowledge and retaining old knowledge by merging the new and past models. One line of research focuses on model averaging, which interpolates between the weights of different LLMs (Wortsman et al., 2022a; b; Eeckt et al., 2022; Yadav et al., 2024; Lin et al., 2023; 2024a). Another line of research relies on the observation that task-specific knowledge largely resides in a subspace of the weight space (Ilharco et al., 2023; Panigrahi et al., 2023; Gueta et al., 2023; Zhu et al., 2024; He et al., 2024), and leverage task vectors or task localization to preserve pre-training knowledge in the fine-tuned models (Panigrahi et al., 2023; Yadav et al., 2024; Yu et al., 2024a).
Architecture-based methods. These methods modify the model’s architecture in training. LoRA (Hu et al., 2022), as the most popular parameter-efficient fine-tuning (PEFT) method, freezes the pre-training weights and introduces low-rank trainable matrices. Variants of LoRA are applied in continual learning for LLMs (Ren et al., 2024; Wang et al., 2023a). However, LoRA is observed to forget less but also learn less than default fine-tuning (Biderman et al., 2024). Apart from LoRA, Adapters (Houlsby et al., 2019) and BitFit (Zaken et al., 2021) are also well-known PEFT methods. In Appendix E.4, we include them as baselines for comparison and find that, while they are slightly less effective than MoFO in mitigating forgetting, their performance on the fine-tuning task is substantially worse than that of MoFO.
Other approaches adaptively expand model capacity or isolate partial weights to mitigate interference between new and old tasks (Wang et al., 2023a; Razdaibiedina et al., 2023). In contrast, MoFO updates a subset of parameters at each iteration, but does not alter the total trainable parameters.
5.2 Block Coordinate Descent
Block Coordinate Descent (BCD) involves iteratively optimizing over a block of coordinates while holding the others constant. The foundational work of Tseng (2001) provides a comprehensive analysis of the convergence properties of BCD under certain conditions. Subsequent research has explored various BCD variants (Hong et al., 2017), including random BCD (Nesterov, 2012; Richtárik & Takáč, 2014; Lu & Xiao, 2015), cyclic BCD (Sun & Hong, 2015; Razaviyayn et al., 2013), and greedy BCD (Nutini et al., 2015). Among these, the greedy variant, also known as Gauss-Southwell BCD method, has drawn attention due to its ability to prioritize coordinates that yield the most substantial improvement in each iteration, thereby potentially accelerating convergence.
In the realm of machine learning, BCD has also found applications (Nutini et al., 2022). For example, Luo et al. (2024) leverages BCD to perform memory-efficient fine-tuning of LLM and Xu & Zhang (2024) uses random masking to perform this. In federated learning, Rothchild et al. (2020) adopts top-$k$ momentum value unsketch rather than our top-$k$ momentum filtering to tackle communication bottleneck and convergence issues. In LLMs, some concurrent works propose BCD-based algorithms leveraging task vectors to enhance fine-tuning performance (Li et al., 2024) and mitigate catastrophic forgetting in multi-task learning (Panda et al., 2024). Our approach can be regarded as a type of greedy BCD adapted to Adam, achieving good performance in fine-tuning tasks and alleviating forgetting.
@@ 6 Conclusion and Limitations
Acknowledgement
The authors would like to express our sincere gratitude to the reviewers for their insightful feedback during the discussion phase. The authors also thank Congliang Chen and Ziniu Li for their helpful suggestions. This paper is supported by NSFC (No. 12326608 and No. 12401409); Hetao Shenzhen-Hong Kong Science and Technology Innovation Cooperation Zone Project (No.HZQSWS-KCCYB-2024016); University Development Fund UDF01001491, the Chinese University of Hong Kong, Shenzhen; Guangdong Provincial Key Laboratory of Mathematical Foundations for Artificial Intelligence (2023B1212010001).
Appendix A Theoretical Analysis
Appendix A is organized into three self-contained parts. A roadmap is provided to guide the reader in locating and understanding each theoretical result efficiently.
Roadmap to Appendix A (Quick Reference).
Appendix A.1: Supplementary Analysis on the Top-$\alpha$ Filter
This section formalizes the top-$\alpha$ filter $\mathrm{FLT}_{\alpha}(\cdot)$ and the induced quantity $\|z\odot\mathrm{FLT}_{\alpha}(z)\|_{1}$ that will serve as our working norm in the analysis. Proposition 1 verifies that this is indeed a norm. Lemma 1 establishes a Lipschitz stability property for filtered outputs, which later lets us pass from the momentum-defined filtering to the gradient-defined filtering. Lemma 2 relates the $L_{1,\text{top-}\alpha}$ norm to standard $L_{p}$ norms, which is used at the very end to translate the convergence bound to any $p\in[1,\infty]$.
Appendix A.2: Proof of Theorem 1 (Convergence of MoFO)
For a high-level narrative, see the proof sketch accompanying Theorem 1 in the main text; here we provide a quick-reference map to the technical steps.
Lemma 3 bounds each step size and the $\ell_{2}$-movement of parameters using the number of active coordinates. Lemma 4 controls the drift of gradients across iterations via $L$-smoothness. Lemma 5 lower-bounds the per-coordinate alignment term $g_{i,t}\hat{m}_{i,t}/\sqrt{\hat{v}_{i,t}}$, which is the driver in the descent inequality. Lemma 6 shows that the bias-corrected momentum $\hat{m}_{t}$ tracks $g_{t}$ in $\ell_{1}$ at rate $O(t^{-1/2})$. Combining these with the descent lemma yields a key inequality in which the filter is $\texttt{FLT}_{\alpha}(m_{t})$:
$\frac{C_{1}}{\sqrt{t}}\|g_{t}\odot\texttt{FLT}_{\alpha}(m_{t})\|_{1}\leq\mathcal{L}(\theta_{t-1})-\mathcal{L}(\theta_{t})+\frac{C_{2}}{t}.$
Finally, Lemma 1 help convert it to the desired inequality with the gradient filter $\texttt{FLT}_{\alpha}(g_{t})$; summing over $t$ gives $\min_{0\leq t\leq T-1}\|\nabla\mathcal{L}(\theta_{t})\odot\texttt{FLT}_{\alpha}(\nabla\mathcal{L}(\theta_{t}))\|_{1}=O(\log T/\sqrt{T})$. By Lemma 2, this implies the same $O(\log T/\sqrt{T})$ rate for any $L_{p}$ norm $\min_{0\leq t\leq T-1}\|\nabla\mathcal{L}(\theta_{t})\|_{p}$.
Appendix A.3: Proof of Theorem 2 (Illustrative example: forgetting mitigation of MoFO)
In an illustrative example with updating ratio $\alpha=1/d$, MoFO converges to a single-attractor minimum and is strictly closer to the pre-training state than Adam, thus attaining a lower pre-training loss. We prove this by mathematical induction.
Appendix A.4: Challenges and Potential Extensions of Theorem 1 to Nonsmooth Objectives
This subsection explains why extending the convergence analysis of Theorem 1 to nonsmooth objectives is nontrivial and outlines several possible future directions.
A.1 Supplementary Analysis on the Top-$\alpha$ Filter